OpenStudy (anonymous):

Glad this site still exists! I have a question on probability that is bothering me. In a promotional effort: customers are encouraged to enter a sweepstakes. Each customer picks 9 numbers between 1-50, inclusive. At the end of the promo period, 9 numbers from 1-50 are drawn, inclusive, and if a customer has 9 numbers matching(without order), the customer wins $5,000,000. What is the probability of the customer winning 5 million dollars? There are other parts to this question but Id rather try myself.

OpenStudy (anonymous):

Oh I want to mention that I tried (50!/9!) but that didnt seem right

OpenStudy (anonymous):

there are \(\binom{50}{9}\) ways to pick the 9 numbers

OpenStudy (anonymous):

if order doesn't matter

OpenStudy (anonymous):

Ohh right the choose thing so 50! ---- 9!(50-9)!

OpenStudy (anonymous):

it would be a miracle if \(\frac{50!}{9!}\) was a probability since that number is certainly not between 0 and 1

OpenStudy (anonymous):

the denominator is \[\binom{50}{9}\] now you have picked a right number up to order there are \(9!\) ways to rearrange the 9 correct numbers

OpenStudy (anonymous):

Ahh ok lol I see so for expected value if I am correct its just this \[\left(\begin{matrix}50 \\ 9\end{matrix}\right)(5,000,000)\]

OpenStudy (anonymous):

Just making sure I am on the right track for this problem

OpenStudy (anonymous):

Thanks for the help Satellite73

OpenStudy (kropot72):

The expected value is given by: \[\large E[X]=\frac{5,000,000}{\left(\begin{matrix}50 \\ 9\end{matrix}\right)}\]

OpenStudy (anonymous):

Hmmm why division?

OpenStudy (anonymous):

Oh wait never mind

OpenStudy (kropot72):

Because the probability of winning is given by: \[\large P(winning)=\frac{1}{\left(\begin{matrix}50 \\ 9\end{matrix}\right)}\]

OpenStudy (anonymous):

Because it has a probability of \[\frac{ 1 }{ 2505433700 }\]

OpenStudy (anonymous):

hold the phonen

OpenStudy (anonymous):

the probability of winning is not \[\frac{1}{\binom{50}{9}}\]

OpenStudy (anonymous):

because you don't have to get the numbers in the exact order

OpenStudy (anonymous):

hmm maybe i am wrong

OpenStudy (anonymous):

But the property of probability has to between 0 and 1

OpenStudy (kropot72):

But the symbol being used is for combinations, not permutations.

OpenStudy (anonymous):

Therefore we cannot leave it at just \[\left(\begin{matrix}50 \\ 9\end{matrix}\right)\]

OpenStudy (anonymous):

no i was thinking that the numerator is \(9!\) the number of ways you can arrange the 9 numbers but maybe i am mistaken

OpenStudy (anonymous):

nvm i am wrong

OpenStudy (anonymous):

Its 50 numbers and you choose 9

OpenStudy (kropot72):

So putting it another way \[\large 50C9=\left(\begin{matrix}50 \\ 9\end{matrix}\right)\] where order does not matter.

OpenStudy (anonymous):

yeah you are right sorry

OpenStudy (anonymous):

Lol its ok thanks for the help :)

OpenStudy (anonymous):

We all have our brain farts

OpenStudy (anonymous):

\[\frac{1}{\binom{50}{9}}\] is right

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

Glad to see this site doing well and thanks dropot72 and satalite73 for your help

OpenStudy (kropot72):

You're very welcome :)

OpenStudy (anonymous):

Keep up the good work